OpenStudy (anonymous):

solve:4x-3(x+5)+x=-9

OpenStudy (anonymous):

4x-3x-15+x=-9 2x=6 x=3

OpenStudy (anonymous):

4x-(3x+15)+x=-9 4x-3x-15+x=-9 2x-15=-9 2x=6 x=3